How many genotypes are possible if a character is controlled by three alleles?

Grade:12

3 Answers

SAHIL
7 years ago
HI KRISHNA PRIYA
WITH THREE ALLELES THERE WILL BE 6 POSSIBLE GENOTYPES THERE. TO GET FREE NOTES REGARDING GENETICS VISIT
THANKYOU
KEEP POSTING YOUR DOUNBTS:}
krishna priya
7 years ago
what is the formula used 
------------------------------------------------------------------------------------------------------------------------
Umakant biswal
7 years ago
@ krishna 
no of possible genotype can be calculated by the following formula 
n(n+1)/2 
where n is the no of alleles 
no of genotype = 3(4)/2 
= 12/2 = 6 
so, in total the no of genotype will be 6 
HOPE IT HELPS 
ALL THE BEST ..